OpenStudy (littlebird):

lim x-> pi/2+ (4/x)secx

geerky42 (geerky42):

Because denominator doesn't approach to 0, you can apply quoient rule. Then try to evaluate \(\lim_{x\to\frac\pi2^+}~ \sec x\)

OpenStudy (littlebird):

Is the quotient rule where you divide by limits? If so, would I then write lim4/limx *limsecx

OpenStudy (littlebird):

I think secx = 1/cosx, and cos(pi/2) = 0, but I then I'm stuck

geerky42 (geerky42):

Yeah. I was more look of \(\dfrac4x\sec x = \dfrac{4\sec x}x\), so \(\lim\dfrac{4\sec x}x = \dfrac{\lim 4\sec x}{\lim x}\) Try imagine the graph of sec x, what happen to sec x as x approaches \(\pi/2\) from right side?

OpenStudy (littlebird):

It approaches -infinity.

geerky42 (geerky42):

Right. and denominator will be just positive value So anything positive multiplied by negative infinity is just negative infinity. Is that clear?

geerky42 (geerky42):

Answer is \(\boxed{-\infty}\)

OpenStudy (littlebird):

Thanks!